OpenStudy (anonymous):

How to convert 0.52 rad into pi form?

OpenStudy (anonymous):

the answer given is pi/6.but ive got no idea how to convert it.

OpenStudy (chaise):

pi/6 It's a good idea to remember the approximations of the basic operations. 45 degrees x (pi/180) pi/4 x 180/pi = 45 Deg --> Rad = pi/180 Rad --> Rad = 180/pi

OpenStudy (anonymous):

but the decimal is in radians.i have no idea how it became pi/6.

OpenStudy (chaise):

pi/6 = 0.52 3.14/6 = 0.52 It's good to memorise some of them. Eg: pi/2 = 1.57 pi/3 = 1.047 pi/4 = 0.785 pi/6 = 0.52 They arn't hard. Just remember per is approx = 3.14

OpenStudy (anonymous):

oh so you just did memory work?no formulas involved?

OpenStudy (chaise):

I told you; Deg --> Rad = x(pi/180) Rad --> Rad = x(180/pi) Eg: 45 x (pi/180) = 0.785 = pi/4 pi/4 x (180/pi) = 45

OpenStudy (anonymous):

Oh ok.i think i got it.thanks for the help :)

OpenStudy (chaise):

My pleasure :)